10. An electric power station that operates at 50 kV and uses a 10:1 step-up ideal transformer is producing 230 MW (Mega-Watt) of power that is to be sent to a big city which is located 120 km away with only 4% loss. Each of the two wires are made of copper (resistivity 1.68 x 10 m). What is the diameter of the wires? cm
Answer #1

voltage at which power is provided to the city=50*10=500 kV

power to be sent=230 MW

current=power/voltage=230*10^6/(500*10^3)=460 A

total loss=0.04*230 MW=9.2*10^6 W

then resistance required=loss/current^2

=43.478 ohms

total length=120+120=240 km=240000 m

as resistance=resistivity*length/area

==>43.478=1.68*10^(-8)*240000/area

==>area=1.68*10^(-8)*240000/43.478=9.2737*10^(-5) m^2

==>pi*(diameter/2)^2=9.2737*10^(-5)

==>diameter=sqrt(4*9.2737*10^(-5)/pi)=0.010866 m

=1.0866 cm